kenzie126:

Solve for x: −5|x + 1| = 10 A- x = 0 B- x = −3 and x = 1 C- x = −1 and x = 3 D- No solution

umm:

So, would be given as\[−5(|x+1|)=10\]Divide -5 from both sides \[\frac{ -5(|x+1|) }{ -5 }=\frac{ 10 }{ -5 }\]\[|x+1|=−2\]Solve absolute value\[|x+1|=−2\] Leaving you with answer choice D because absolute value cannot be less than 0.